# To be or not to be, but where?

### v1

@@ 8.3 Discussion
Within this approach, we have directly connected insights from quantum gravity, and its lack of gauge-invariant local observables, to the measurement problem in quantum mechanics. Although much work needs to be done, if this program succeeds, it would mean that all classical measurements are interconnected by a web of measurement dualities. The discussion was limited to bipartite systems, but typically we also consider an observer in a quantum measurement, which nowadays is much better discussed and formalized in the context of quantum reference frames [^31] There is a large literature on this topic [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, e.g.]. For a recent paper, see [87, and references 1-23 therein].. In this program, changes of TPS appear naturally as one considers transformations between quantum systems to select the one that will be the reference frame, very similar to what we do in special relativity to choose the preferred inertial frame considered at rest. However, to the best of my knowledge, the status of the program remains mostly kinematical, in the sense that they do not consider the change of TPS due to time evolution for the same quantum system, similar to a classical frame having to update their local inertial frame as they move in a dynamical spacetime. Three recent papers that delve into dynamics are [88, 89, 90]. These works, together with the formalism introduced in [91], might allow for an extension of this proposal to infinite-dimensional Hilbert spaces and help better quantify the mathematical assumptions surrounding the existence of a large and continuous set of different TPS related to classical measurements.
@@ 9 We are inside the box!
The most vivid image you can retrieve from this whole story is that we are inside that mysterious box! In fact, due to the seemingly unavoidable non-local nature of gauge-invariant observables in quantum gravity, to some extent, we are the box. This is why holography has been one of the main entries into quantum gravity over the last few decades.
The issue we highlight here is that, generally, we are not using holography to solve elementary problems in physics involving gravity and quantum phenomena [^32] As usual, there are counterexamples such as [92, 93, and references therein]. Typically, we rely on perturbative quantum gravity under standard canonical quantization. Despite preventing us from establishing gauge-invariant local observables, we write down models in these contexts as if that were not an issue. And, in truth, it works—to the extent that we can conduct experiments in our labs and explain them successfully. Naturally, this raises two questions: why does it work, and where should we expect it not to work?
In [71], an answer to the first question is proposed. For classical gravity, they demonstrate that certain classical configurations can be gravitationally dressed in a way that the dressing depends only on the Poincaré charges (energy, momentum, and angular momentum) of the configuration. Different configurations with the same Poincaré charges are indistinguishable from outside their region, suggesting a way to localize information classically. In the quantum context, they argue that perturbative quantum gravity in the presence of global symmetries allows for the definition of localized information or “gravitational qubits,” which can be localized to regions in such a way that their information content is not accessible outside the region. Meanwhile, [94] seems to obtain a general and fully non-perturbative framework for constructing dynamical reference frames in generally covariant theories. By constructing relationally local observables throughout spacetime that satisfy microcausality in the bulk, they establish what they call ”relational locality.”
@@ 10 Acknowledgements
This work is the result of numerous discussions with many people over the last two years. I am deeply grateful to the EmerGe (Emergent Geometries) collaboration, which includes Eugenia Colafranceschi, Andrea Di Biagio, Joakim Flinckman, Jan Głowacki, Niels Linnemann, and Florian Niedermann. Collaborating with them has allowed me to deepen my knowledge and interest in classical and quantum subsystems. This interest began during the Sejny Summer Institute in 2021, sparked by conversations with Pierre Martin-Dussaud. I have discussed using changes in the TPS to address the measurement problem with Andrea Di Biagio, Flavio Del Santo, Patrick Fraser, Arsalan Adil, Leon Loveridge, Anne-Catherine de la Hamette, and Jan Głowacki. The last three provided feedback on the discussions presented in this paper. I am also thankful to Vasileios Fragkos for his feedback on the section on GIE experiments, as well as to Benjamin Bose and Florian Niedermann for reviewing parts of the draft. Finally, many ideas concerning the measurement problem emerged from supervising the master thesis of Nicolas Boulles, with whom I had extensive discussions on the topic.

### v2

@@ 8.3 Discussion
Within this approach, we have directly connected insights from quantum gravity and its lack of gauge-invariant local observables to the measurement problem in quantum mechanics. Although much work needs to be done, if this program succeeds, it would mean that all classical measurements are interconnected by a web of measurement dualities. The discussion was limited to bipartite systems, but typically we also consider an observer in a quantum measurement, which nowadays is much better discussed and formalized in the context of quantum reference frames [^36] There is a large literature on this topic [88, 89, 90, 91, 92, 93, 94, 95, 96, 87, 97, 98, e.g.]. For a recent paper, see [99, and references 1-23 therein].. In this program, changes of TPS appear naturally as one considers transformations between quantum systems to select the one that will be the reference frame, very similar to what we do in special relativity to choose the preferred inertial frame considered at rest. However, the status of the program remains mostly kinematical, in the sense that they do not consider the change of TPS due to time evolution for the same quantum system, similar to a classical frame having to update its local inertial frame as it moves in a dynamical spacetime. Three recent papers that delve into dynamics are [100, 101, 102]. These works, together with the formalism introduced in [103], might allow for an extension of this proposal to infinite-dimensional Hilbert spaces and help better quantify the mathematical assumptions surrounding the existence of a large and continuous set of different TPS related to classical measurements.
@@ 9 We are inside the box!
The most vivid image you can retrieve from this whole story is that we are inside that initial mysterious box! In fact, due to the seemingly unavoidable non-local nature of gauge-invariant observables in quantum gravity, to some extent, we are the box. This is why holography has been one of the main entries into quantum gravity over the last few decades.
The issue we highlight here is that, generally, we are not using holography to solve elementary problems in physics involving gravity and quantum phenomena [^37] As usual, there are counterexamples such as [104, 105, and references therein].. Typically, we rely on perturbative quantum gravity under standard covariant quantization. Despite preventing us from establishing gauge-invariant local observables, we write down models in these contexts as if that were not an issue. And, in truth, it works—to the extent that we can conduct experiments in our labs and explain them successfully. Naturally, this raises two questions: why does it work, and where should we expect it not to work?
In [78], an answer to the first question is proposed. For classical gravity, they demonstrate that certain classical configurations can be gravitationally dressed in a way that the dressing depends only on the Poincaré charges (energy, momentum, and angular momentum) of the configuration. Different configurations with the same Poincaré charges are indistinguishable from outside their region, suggesting a way to localize information classically. In the quantum context, they argue that perturbative quantum gravity in the presence of global symmetries allows for the definition of localized information or “gravitational qubits,” which can be localized to regions in such a way that their information content is not accessible outside the region. Meanwhile, [106] seems to obtain a general and fully non-perturbative framework for constructing dynamical reference frames in generally covariant theories. By constructing relationally local observables throughout spacetime that satisfy microcausality in the bulk, they establish what they call “relational locality.”
@@ 10 Acknowledgements
This work is the result of numerous discussions with many people over the last two years. I am deeply grateful to the EmerGe (Emergent Geometries) collaboration, which includes Eugenia Colafranceschi, Andrea Di Biagio, Joakim Flinckman, Jan Głowacki, Niels Linnemann, and Florian Niedermann. Collaborating with them has allowed me to deepen my knowledge and interest in classical and quantum subsystems. This interest began during the Sejny Summer Institute in 2021, sparked by conversations with Pierre Martin-Dussaud. I have discussed using changes in the TPS to address the measurement problem with Andrea Di Biagio, Flavio Del Santo, Patrick Fraser, Arsalan Adil, Leon Loveridge, Anne-Catherine de la Hamette, Sören Holst, and Jan Głowacki. The last three provided feedback on the discussions presented in this paper. I am also thankful to Vasileios Fragkos for his feedback on the section on GIE experiments, as well as to Benjamin Bose and Florian Niedermann for reviewing parts of the draft. I am also thankful to Emily Adlam, Tales Rick Perche and Shadi Ali Ahmad for careful reading and further comments on the draft. Finally, many ideas concerning the measurement problem emerged from supervising the master’s thesis of Nicolas Boulles, with whom I had extensive discussions on the topic. This work was supported by the Swedish Research Council (grant number 2022-01893VR).